OpenStudy (anonymous):

A rectangle is 5 times as long as it is wide. It's perimeter is 180 feet. What is the area of the rectangle, in square feet?

OpenStudy (eyust707):

We need to write down two equations

OpenStudy (eyust707):

"A rectangle is 5 times as long as it is wide" Can you think of an equation for this statement?

OpenStudy (anonymous):

5*lw?

OpenStudy (eyust707):

gotta have an = sign in there somewhere...

OpenStudy (eyust707):

"5w is the same as l"

OpenStudy (eyust707):

How about that one?

OpenStudy (anonymous):

5w?

OpenStudy (eyust707):

you need an = sign!

OpenStudy (anonymous):

5=w?

OpenStudy (eyust707):

no no w is not 5.. we need to solve for w

OpenStudy (anonymous):

sorry.

OpenStudy (eyust707):

"5w is the same as l" 5w = l

OpenStudy (eyust707):

okay so one of our equations is 5w = l

OpenStudy (anonymous):

ok....

OpenStudy (eyust707):

How do you find the perimeter?

OpenStudy (anonymous):

add the sides together

OpenStudy (eyust707):

sweet! can you write that using "w" and "l"

OpenStudy (anonymous):

5+1+5+1=12 180/12=15 5*15=75 75*15=1125